An external magnetic field changes as a function of time such that: (1) it is momentarily zero at time "t=0" , and (2) it produces an electric field "E=x\hat{y}-y\hat{z}" that is constant in time. Determine the magnitude and direction of the magnetic field as a function of time


1
Expert's answer
2021-05-06T17:23:07-0400

Given,

Electric field is given as "(E)= x\hat{y}-y\hat{z}"

We know that,

"\Rightarrow \nabla \times E=\frac{-dB}{dt}"


"\Rightarrow [\frac{\partial }{\partial x}\hat{i}+\frac{\partial }{\partial y}\hat{j}+\frac{\partial }{\partial x}\hat{z}]\times[x\hat{y}-y\hat{z}]=-\frac{dB}{dt}"


"\Rightarrow \hat{i}+\hat{k}=\frac{-dB}{dt}"

"B= -t(\hat{i}+\hat{k})"